# Auto-Generating Earnings Report Analysis via an Augmented LLM

### v1

@@ 4.3 Quantitative comparison
Model Correctness (1: worst, 10: best) Semantic Similarity (smaller is better)
@@ 4.4 Qualitative comparison
Table 4: Qualitative comparison of the examples of earnings analysis for our augmented LLM, Llama-2-7b and GPT-3.5.
Our Financially-augmented LLM
Question:
What is Broadcom Inc.?
Generated Answer:
@@
Broadcom Inc. is a technology company that designs, develops, and supplies semiconductor and infrastructure software solutions.
Question:
What is the company’s quarterly dividend for the third quarter of Fiscal Year 2023?
Generated Answer:
4.60
Ground-truth Answer:
The company’s quarterly dividend for the third quarter of Fiscal Year 2023 is \$4.60 per share.
Llama-2-7b
Question:
What is Broadcom Inc.?
Generated Answer:
@@
1. An American manufacturer of semiconductor devices and infrastructure software solutions.
2. A software company that provides solutions for managing applications.
Question:
What is the company’s quarterly dividend for the third quarter of Fiscal Year 2023?
Generated Answer:
368.40
GPT-3.5
Question:
What is Broadcom Inc.?
Generated Answer:
@@
Broadcom Inc. is a global technology leader that designs, develops, and supplies semiconductor and infrastructure software solutions.
Question:
@@
What is the company’s quarterly dividend for the third quarter of Fiscal Year 2023?
Generated Answer:
The company’s quarterly dividend for the third quarter of Fiscal Year 2023 is \$4.60.
@@
We conducted a rigorous qualitative comparison of our augmented model with two baseline models. To ensure a fair evaluation, all three models were provided with the same quarterly report document, and we employed an identical embedding model for vector indexing. Table 4 showcases an illustrative example of a question pertaining to earnings analysis, along with the generated answers from each model. The qualitative analysis presented in the table distinctly demonstrates that our Financially-augmented model consistently produces answers that are notably contextually relevant to the given question when compared to Llama-2, and similar to GPT-3.5 in this regard. This empirical evidence underscores the efficacy and superior performance of our model within the specialized domain of financial analysis.

### v3

@@ 5.3 Quantitative comparison
Model
Correctness
(1: worst, 10: best)
Semantic Similarity
(smaller is better)
@@ 5.4 Qualitative comparison
Table 4: Qualitative comparison of model responses across different financial question types.
(a) Question Type: General Company Information
Question: What is Broadcom Inc.?
Model Generated Answer
@@
Our Financially-augmented LLM Broadcom Inc. is a technology company that designs, develops, and supplies semiconductor and infrastructure software solutions.
@@
Llama-2-7b 1. An American manufacturer of semiconductor devices and infrastructure software solutions. 2. A software company that provides solutions for managing applications.
@@
GPT-3.5 Broadcom Inc. is a global technology leader that designs, develops, and supplies semiconductor and infrastructure software solutions.
(b) Question Type: Financial Metric Extraction Query
@@
Question: What is the company’s quarterly dividend for the third quarter of Fiscal Year 2023?
Ground-truth: \$4.60 per share.
Model Generated Answer
Our Financially-augmented LLM \$4.60
Llama-2-7b \$368.40
GPT-3.5 The company’s quarterly dividend for the third quarter of Fiscal Year 2023 is \$4.60.
@@
We conducted a rigorous qualitative comparison of our augmented model with two baseline models. To ensure a fair evaluation, all three models were provided with the same quarterly report document, and we employed an identical embedding model for vector indexing. Table 4 presents examples of questions related to earnings analysis, along with the corresponding generated answers from each model. The qualitative analysis reveals clear distinctions in the quality and contextual relevance of responses.
In the general company information question, all three models, the open-source general-purpose Llama-2-7b, our Financially-augmented LLM, and the commercial GPT-3.5, produce correct and sufficiently detailed descriptions of Broadcom Inc. This indicates that broad corporate identity queries can be reliably answered even without financial-domain specialization. However, in the financial metric extraction query, our model provides a more accurate and semantically aligned answer than the open-source baseline and shows a level of correctness that closely matches the response from the powerful commercial GPT-3.5 model.
Our Financially-Augmented LLM.
Our model consistently demonstrates high contextual accuracy and precise numerical reasoning. For instance, in the dividend question, the generated response (”4.60”) directly matches the ground-truth value and preserves factual consistency with the source document. The model effectively retrieves and integrates the relevant context from the financial report, yielding concise and unambiguous answers. This indicates that the retrieval-augmented instruction tuning process allows the model to better internalize document structure and financial semantics. Moreover, it produces domain-appropriate terminology (e.g., “semiconductor and infrastructure software solutions”) comparable in style and factual fidelity to human financial analysts.
Llama-2-7b Baseline.
The baseline Llama-2-7b model exhibits limited grounding and weaker numerical reasoning. Although it provides a generally correct company description, the responses show redundancy and lack of focus, as evidenced by the listing of multiple partial statements. For numerical queries, the model frequently produces spurious values (e.g., ”368.40”), demonstrating its inability to extract and reason over financial figures from context. These errors highlight the lack of document-grounded alignment and domain-specific adaptation in the base LLM.
GPT-3.5 Teacher Model.
GPT-3.5 generates highly coherent and semantically accurate responses, often matching or exceeding the clarity of human-written financial summaries. Its answer for the dividend question exactly reproduces the ground truth, confirming strong factual precision and contextual comprehension. However, GPT-3.5 relies heavily on its large-scale pretrained knowledge and less on explicit document retrieval, occasionally leading to generic phrasing or assumptions when context is ambiguous. Despite this, it serves as an effective teacher model for generating context-grounded instruction data due to its consistent reasoning quality.
Overall Comparison.
Overall, our Financially-Augmented LLM achieves a balance between GPT-3.5’s reasoning depth and Llama-2’s fine-tuning adaptability. The model effectively integrates retrieval context to ensure factual correctness while maintaining natural linguistic fluency. Qualitatively, it shows a clear improvement in interpretability, numeric reliability, and contextual grounding compared to the baseline. This result substantiates the effectiveness of retrieval-augmented instruction tuning in enabling domain-specialized financial reasoning within smaller, efficient LLMs.
5.5 More Success and Failure Cases
We provide additional success and failure cases of our model across different question types in Table 5, including different key financial indicators including dividend queries, revenue-related metrics, period identification, and segment-level breakdowns. These failure cases span multiple companies (Broadcom, NVIDIA, and AMD), illustrating that errors may occur not only in numeric extraction but also in semantic alignment with the ground-truth context. For each example, we list the question, the financial metric type, the model’s generated answer, and the judgment from the GPT-4 judge model.
Table 5: Representative success and failure cases of our model across various financial metric types.
Question Metric Type Generated Answer Judge Model Reasoning Score
What is the company’s quarterly dividend for the third quarter of Fiscal Year 2023? dividend The company did not declare or pay any dividends for the third quarter of Fiscal Year 2023. The model states that no dividend was declared, contradicting the ground truth that a dividend was paid. 1
What is the reported revenue for the third quarter of Fiscal Year 2023? revenue-related The reported revenue for the third quarter of Fiscal Year 2023 is \$13,507 million, up 101% from the same period last year. The model provides revenue of \$13,507M, while the correct figure is \$8,876M, indicating a significant numeric error. 5
Can you provide a breakdown of the revenue by business segment? segment-level breakdowns Based on the provided context, AMD’s revenue by business segment for the three months ended July 1, 2023 and June 25, 2022 are as follows: * Data Center: \$1,321 million in net revenue for Q2 2023 and \$1,486 million in Q2 2022. * Client: \$998 million in net revenue for Q2 2023 and \$2,152 million in Q2 2022. * Gaming: \$1,581 million in net revenue for Q2 2023 and \$1,655 million in Q2 2022. * Embedded: \$1,459 million in net revenue for Q2 2023 and \$1,257 million in Q2 2022. Note that these figures are based on management’s judgment and allocation of expenses to the reportable segments, and may not necessarily reflect actual profitability by segment. The model fabricates a detailed AMD revenue breakdown even though the reference document does not provide such information. 1
What is the quarterly period ended date? period identification Based on the context provided, the quarterly period ended date is July 1, 2023. The model answers July 1, 2023, but the correct period end date is July 30, 2023. 7
What is the revenue of NVIDIA for second quarter fiscal year 2024? revenue-related NVIDIA’s revenue for the second quarter of fiscal year 2024 was \$13.51 billion, up 101% from the same period last year and up 88% sequentially. The model correctly provides \$13.51B and correctly add the percentage increase. 10
6 Related Work
Large Language Models (LLMs) have been increasingly applied to financial analysis, with early domain-specific models like BloombergGPT [15] and FinGPT [16] demonstrating the value of financial adaptation. Recent studies extend this trend toward more sophisticated analytical tasks. Jadhav and Mirza [7] provide a comprehensive review of 84 studies on LLMs in equity markets, emphasizing their applications in forecasting and portfolio management. Regulatory perspectives, such as the ESMA–Alan Turing Institute report [3], further highlight the importance of responsible and verifiable adoption in finance.
Instruction tuning has emerged as an efficient strategy for aligning LLMs with domain-specific objectives. Han et al. [4] present a comprehensive survey of instruction tuning methods, emphasizing alignment-centric paradigms. This paper also shows that data quality and domain diversity critically influence downstream task performance. In finance, Hirano and Imajo [5] propose building instruction-tuned LLMs without explicit instruction datasets via continual pretraining and model merging, and Tanabe et al. [10] introduce JaFIn, a Japanese financial instruction dataset that enhances domain adaptability.
Retrieval-Augmented Generation (RAG) further improves factual grounding and contextual accuracy in financial reasoning. Zeng [17] presents QuantMCP, a retrieval-grounded framework linking LLMs to verified financial data sources, while Wang et al. [14] evaluate RAG and fine-tuning approaches for financial-statement analysis, highlighting their complementary strengths. Despite these advances, the integration of retrieval augmentation and instruction tuning for automated earnings report analysis remains unexplored. This work introduces RAG-IT, a framework unifying retrieval-augmented instruction data generation and fine-tuning to address this gap in financial LLM research.
@@ 7 Conclusion
Nevertheless, our approach has several limitations. RAG-IT may struggle with highly nuanced questions that require deep domain expertise, multi-hop reasoning over various financial metrics, or precise numerical computation. Its accuracy also depends heavily on the completeness and quality of retrieved evidences; missing or poorly retrieved context can degrade performance. Inference requires running both retrieval and generation, which introduces additional latency and computational cost compared to standard LLM inference. These limitations highlight important directions for future research, including more robust retrieval methods, improved reasoning capabilities, and more efficient architectures for financial analysis.